OpenStudy (anonymous):

The sum of twice a number and another number is 24. The difference of twice the first number and the other number is 12. Which system would model this situation, and what is the solution?

OpenStudy (midhun.madhu1987):

Let the numbers be x and y. 2x + y = 24 2x - y = 12 Solve them...

OpenStudy (anonymous):

lol but how im not very good in math

OpenStudy (midhun.madhu1987):

Add those 2 equations.... and y will be eliminated...find x.

OpenStudy (anonymous):

A. 2(x + y) = 24 2(x – y) = 12 Solution: (9, 6)